# Compatibility of canonical $\ell$-adic local systems on Shimura varieties

### v1

@@ 2. Recollection on companions
classifying continuous semisimple $\lambda$-adic representations whose Frobenius eigenvalues are all plain of characteristic $p$ [^3] We give a few details not made so explicit in [7, §6]. Either the Weil number or plainness conclusions can be checked after replacing $X$ by a finite étale cover. Thus for a given representation $\hat{\Pi}_{\lambda}^{\mathrm{red}}\twoheadrightarrow H$ (which in the two cases we would assume factors through $\hat{\Pi}_{\lambda}^{\mathrm{plain}}$ or $\hat{\Pi}_{\lambda}^{\mathrm{mot}}$), we can assume it factors through $H^{\mathrm{geo},0}\cdot Z_{H^{0}}^{0}$, with $H^{\mathrm{geo}}$ denoting the geometric monodromy group. Projecting, we get a representation $\hat{\Pi}^{\mathrm{red}}_{\lambda}\to H^{\mathrm{geo},0}/H^{\mathrm{geo},0}\cap Z_{H^{0}}^{0}$ that factors, by semisimplicity of geometric monodromy, through $\hat{\Pi}_{\lambda}$; again replacing $X$ by a finite étale cover, we may assume this lifts to a homomorphism $\hat{\Pi}_{\lambda}\to H^{\mathrm{geo},0}$. Similarly projecting to $Z^{0}_{H^{0}}/H^{\mathrm{geo},0}\cap Z_{H^{0}}^{0}$ and replacing $X$ by a finite étale cover that allows us to lift to $Z^{0}_{H^{0}}$, we get a homomorphism $\hat{\Pi}_{\lambda}^{\mathrm{red}}\to Z^{0}_{H^{0}}$ that factors through $\mathrm{Hom}(A,\mathbb{G}_{m})$, with $A=\overline{\mathbb{Z}}_{\lambda}^{\times}$, $\mathcal{W}_{p}$, or $\mathcal{P}_{p}$ according to our assumptions on the original representation. (Note that $\ker(\hat{\Pi}_{\lambda}^{\mathrm{red}}\to\mathrm{Hom}(A,\mathbb{G}_{m}))$ is isomorphic to $\ker(\hat{\Pi}_{\lambda}\to\hat{\mathbb{Z}})$ and factors through $H^{\mathrm{geo},0}$ in our $H$-representation.) In summary, after replacing $X$ by a finite cover, we can assume our homomorphism is a product of a $\hat{\Pi}_{\lambda}\to H^{\mathrm{geo},0}$ and a $\mathrm{Hom}(A,\mathbb{G}_{m})\to Z^{0}_{H^{0}}\cong\mathbb{G}_{m}^{r}$, i.e. an $r$-tuple $(\alpha_{1},\ldots,\alpha_{r})\in A^{r}$. In any finite-dimensional representation of $H$, the former sends any $\mathrm{Frob}_{x}$ to a matrix with pure weight 0 eigenvalues, and the latter sends $\mathrm{Frob}_{x}$ to a product $\prod_{i=1}^{r}\alpha_{i}^{d_{i}\deg(x)}$, where the $d_{i}$ are integers depending on the representation and $\deg(x)=[\kappa(x):\mathbb{F}_{p}]$. Thus representations of $\hat{\Pi}_{\lambda}^{\mathrm{mot}}$ and $\hat{\Pi}_{\lambda}^{\mathrm{plain}}$ give $\lambda$-adic representations with Weil number and plain eigenvalues as claimed. The converse is similar. $\mathcal{W}_{p}$ and $\mathcal{P}_{p}$ are both independent of $\lambda$, so using the main theorem [7, Theorem 1.4.1] on $\lambda$-independence of $\hat{\Pi}_{\lambda}$ we can define groups over $\overline{\mathbb{Q}}$, $\hat{\Pi}^{\mathrm{plain}}=\hat{\Pi}\times_{\hat{\mathbb{Z}}}\mathrm{Hom}(\mathcal{P}_{p},\mathbb{G}_{m})$ (and likewise $\hat{\Pi}^{\mathrm{mot}}$, as already done in [7, §6.1]) such that for all, $\lambda$ $\hat{\Pi}^{\mathrm{plain}}\times_{\overline{\mathbb{Q}}}\overline{\mathbb{Q}}_{\lambda}$ is isomorphic, uniquely up to conjugation by $\hat{\Pi}_{\lambda}^{\mathrm{plain},0}$, to $\hat{\Pi}_{\lambda}^{\mathrm{plain}}$.
@@
In particular, in the setting of the Lemma, we have a surjection $\hat{\Pi}_{\lambda}^{\mathrm{red}}\twoheadrightarrow\hat{\Pi}_{\lambda}^{\mathrm{plain}}\twoheadrightarrow H_{\overline{\mathbb{Q}}_{\lambda}}$. Descending the second arrow to $\overline{\mathbb{Q}}$, we get $\hat{\Pi}^{\mathrm{plain}}\twoheadrightarrow H$ (unique up to $H^{0}$-conjugation), which in turn for any $\lambda^{\prime}$ extends to $\hat{\Pi}_{\lambda^{\prime}}^{\mathrm{plain}}\twoheadrightarrow H_{\overline{\mathbb{Q}}_{\lambda^{\prime}}}$, which classifies the companion $\rho_{\lambda\leadsto\lambda^{\prime}}$.

### v2

@@ 2. Recollection on companions
that we claim classifies continuous semisimple $\lambda$-adic representations whose Frobenius eigenvalues are all plain of characteristic $p$. The analogous claim is made in [8, §6] with $\mathcal{W}_{p}$ in place of $\mathcal{P}_{p}$, and in covering both cases and the case of $p$-adic coefficients we will take the opportunity to give a few more details than loc. cit.
For any $\lambda$, define $\hat{\Pi}_{\lambda}^{\mathrm{plain}}$ and $\hat{\Pi}_{\lambda}^{\mathrm{mot}}$ to be the quotients of $\hat{\Pi}_{\lambda}^{\mathrm{red}}$ such that:
  • If $\lambda\nmid p$, a homomorphism $\rho\colon\hat{\Pi}_{\lambda}^{\mathrm{red}}\to H$ factors through $\hat{\Pi}_{\lambda}^{\mathrm{plain}}$ (respectively, $\hat{\Pi}_{\lambda}^{\mathrm{mot}}$) if and only if for all representations $r\colon H\to\mathrm{GL}_{N}$, the $\lambda$-adic representation $r\circ\rho_{\lambda}$ associated to $r\circ\rho$ satisfies: for all closed points $x\in X$, all eigenvalues of $r\circ\rho_{\lambda}(\mathrm{Frob}_{x})$ are plain of characteristic $p$ (respectively, are #$\#\kappa(x)$-Weil numbers).
  • If $\lambda|p$, a homomorphism $\rho\colon\hat{\Pi}_{\lambda}^{\mathrm{red}}\to H$ factors through $\hat{\Pi}_{\lambda}^{\mathrm{plain}}$ (respectively, $\hat{\Pi}_{\lambda}^{\mathrm{mot}}$) if and only if for all representations $r\colon H\to\mathrm{GL}_{N}$, the object $\mathcal{F}$ of $\operatorname{F-Isoc}^{\dagger}(X)$ associated to $r\circ\rho$ satisfies: for all closed points $x\in X$, the eigenvalues of the $\kappa(x)$-linearized Frobenius on $x^{*}\mathcal{F}$ are plain of characteristic $p$ (respectively, are #$\#\kappa(x)$-Weil numbers).
[…]
Lemma 2.2.
For any finite place $\lambda$ of $\overline{\mathbb{Q}}$, there are canonical isomorphisms
$\hat{\Pi}_{\lambda}^{\mathrm{plain}}\to\hat{\Pi}_{\lambda}\times_{\hat{\mathbb{Z}}}\mathrm{Hom}(\mathcal{P}_{p},\mathbb{G}_{m})$
and
$\hat{\Pi}_{\lambda}^{\mathrm{mot}}\to\hat{\Pi}_{\lambda}\times_{\hat{\mathbb{Z}}}\mathrm{Hom}(\mathcal{W}_{p},\mathbb{G}_{m}).$
Proof.
In each case, the first factor of the map is the projection to the pro-semisimple quotient. To define the second factor, we construct the dual map $A\to\mathrm{Hom}(G,\mathbb{G}_{m})$ for $(A,G)$ equal to $(\mathcal{P}_{p},\hat{\Pi}_{\lambda}^{\mathrm{plain}})$ or $(\mathcal{W}_{p},\hat{\Pi}_{\lambda}^{\mathrm{mot}})$ by restricting the source of the maps $\Xi_{\lambda}$.
  • Suppose $\lambda\nmid p$. For $\alpha\in\mathcal{P}_{p}$ (respectively, $\alpha\in\mathcal{W}_{p}$), viewing $\Xi_{\lambda}(\alpha)$ as a homomorphism $\Pi\to\overline{\mathbb{Q}}_{\lambda}^{\times}$, by construction $\Xi_{\lambda}(\alpha)(\mathrm{Frob}_{x})=\alpha^{\#\kappa(x)}$; thus $\Xi_{\lambda}(\alpha)$ factors through $\hat{\Pi}_{\lambda}^{\mathrm{plain}}$ (respectively $\hat{\Pi}_{\lambda}^{\mathrm{mot}}$).
  • Suppose $\lambda|p$. For $\alpha\in\mathcal{P}_{p}$ (respectively, $\alpha\in\mathcal{W}_{p}$), $\Xi_{\lambda}(\alpha)$, viewed as a rank 1 object of $\operatorname{F-Isoc}(\mathrm{Spec}(\mathbb{F}_{p}))\otimes_{\mathbb{Q}_{p}}\overline{\mathbb{Q}}_{\lambda}$, has the property that for all closed points $x$, $x^{*}(\Xi_{\lambda}(\alpha))$ is the rank-1 object of $\operatorname{F-Isoc}(\mathrm{Spec}(\kappa(x)))\otimes_{\mathbb{Q}_{p}}\overline{\mathbb{Q}}_{\lambda}$ whose #$\#\kappa(x)$-linearized Frobenius has eigenvalue $\alpha^{\#\kappa(x)}$; thus $\Xi_{\lambda}(\alpha)$ factors through $\hat{\Pi}_{\lambda}^{\mathrm{plain}}$ (respectively, $\hat{\Pi}_{\lambda}^{\mathrm{mot}}$).
Let $Z^{\mathrm{red},0}$, $Z^{\mathrm{plain},0}$, and $Z^{\mathrm{mot},0}$ be the connected components of the identity of the centers of $\hat{\Pi}_{\lambda}^{\mathrm{red}}$, $\hat{\Pi}_{\lambda}^{\mathrm{plain}}$, and $\hat{\Pi}_{\lambda}^{\mathrm{mot}}$. For any pro-reductive group $G$, we have a short exact sequence
$1\to Z_{G}^{0}\to G\to G^{\mathrm{ss}}\to 1$
with $G^{\mathrm{ss}}$ the maximal pro-semisimple quotient. In all three cases under consideration, $G^{\mathrm{ss}}=\hat{\Pi}_{\lambda}$, so we see that
$\ker(\hat{\Pi}_{\lambda}^{\mathrm{red}}\to\hat{\Pi}_{\lambda}^{\mathrm{mot}})=\ker(Z^{\mathrm{red},0}\to Z^{\mathrm{mot},0})$
and
$\ker(\hat{\Pi}_{\lambda}^{\mathrm{red}}\to\hat{\Pi}_{\lambda}^{\mathrm{plain}})=\ker(Z^{\mathrm{red},0}\to Z^{\mathrm{plain},0}).$
From the first part of the argument of [8, Lemma 3.5.1], we want to show that for each possibility $G=\hat{\Pi}_{\lambda}^{\mathrm{red}}$, $G=\hat{\Pi}_{\lambda}^{\mathrm{plain}}$, $G=\hat{\Pi}_{\lambda}^{\mathrm{mot}}$ and (respectively) $A=\overline{\mathbb{Z}}_{\lambda}^{\times}$ ($\lambda\nmid p$) or ¯$\overline{\mathbb{Q}}_{\lambda}^{\times}$ ($\lambda|p$), $A=\mathcal{P}_{p}$, $A=\mathcal{W}_{p}$, the two maps
[…]
are both isomorphisms. For $G=\hat{\Pi}_{\lambda}^{\mathrm{red}}$ and $\lambda\nmid p$, this is [8, Proposition 3.5.2], and for $G=\hat{\Pi}_{\lambda}^{\mathrm{red}}$ and $\lambda|p$, this is Lemma 2.1. We now explain how the other cases follow formally. First note that the map $f$ is an isomorphism in all three cases. To see this, we observe from the definitions that
¯${\overline{\mathbb{Q}}_{\lambda}^{\times}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Hom}(\hat{\Pi}_{\lambda}^{\mathrm{red}},\mathbb{G}_{m})}$${\mathcal{P}_{p}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Hom}(\hat{\Pi}_{\lambda}^{\mathrm{plain}},\mathbb{G}_{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
is for $\lambda|p$ a pullback square and is moreover modulo torsion a pullback square; the same holds for $\lambda\nmid p$ with ¯$\overline{\mathbb{Z}}_{\lambda}^{\times}$ in place of ¯$\overline{\mathbb{Q}}_{\lambda}^{\times}$, and with $\mathcal{W}_{p}$ and $\hat{\Pi}_{\lambda}^{\mathrm{mot}}$ replacing $\mathcal{P}_{p}$ and $\hat{\Pi}_{\lambda}^{\mathrm{plain}}$. Thus $\mathcal{P}_{p}/\mu_{\infty}\xrightarrow{\sim}\mathrm{Hom}(\hat{\Pi}_{\lambda}^{\mathrm{plain}},\mathbb{G}_{m})/\mathrm{Hom}(\hat{\Pi}_{\lambda}^{\mathrm{plain}},\mathbb{G}_{m})_{\mathrm{tors}}$ for all $\lambda$, and likewise for $\hat{\Pi}_{\lambda}^{\mathrm{mot}}$. The same holds with $X$ replaced by any finite étale connected cover, so $\mathcal{P}_{p}/\mu_{\infty}\xrightarrow{\sim}\mathrm{Hom}((\hat{\Pi}_{\lambda}^{\mathrm{plain}})^{0},\mathbb{G}_{m})$ is also an isomorphism (again, likewise for $\hat{\Pi}_{\lambda}^{\mathrm{mot}}$).
Injectivity of the map $g$ holds for any reductive group. For surjectivity, let $G$ be one of $\hat{\Pi}_{\lambda}^{\mathrm{mot}}$ or $\hat{\Pi}_{\lambda}^{\mathrm{plain}}$. Any element of $\mathrm{Hom}(Z_{G}^{0},\mathbb{G}_{m})\subset\mathrm{Hom}(Z^{\mathrm{red},0},\mathbb{G}_{m})$ is (by Lemma 2.1 and [8, Proposition 3.5.2]) the restriction to $Z^{\mathrm{red},0}$ of $\Xi_{\lambda}(\alpha)$ for some $\alpha\in\overline{\mathbb{Z}}_{\lambda}^{\times}$ ($\lambda\nmid p$) or $\alpha\in\overline{\mathbb{Q}}_{\lambda}^{\times}$ ($\lambda|p$). Since $\ker(\hat{\Pi}_{\lambda}^{\mathrm{red}}\to G)=\ker(Z^{\mathrm{red},0}\to Z_{G}^{0})$, $\Xi_{\lambda}(\alpha)\colon\hat{\Pi}_{\lambda}^{\mathrm{red}}\to\mathbb{G}_{m}$ must factor through $G\to\mathbb{G}_{m}$; this (as above) forces $\alpha$ to lie in $\mathcal{W}_{p}$ or $\mathcal{P}_{p}$ according to the choice of $G$. We conclude that $g\circ f$ is surjective, hence $g\circ f$ is an isomorphism (as are both $g$ and $f$), hence
$\hat{\Pi}_{\lambda}^{\mathrm{plain}}\to\hat{\Pi}_{\lambda}\times_{\hat{\mathbb{Z}}}\mathrm{Hom}(\mathcal{P}_{p},\mathbb{G}_{m})$
and
$\hat{\Pi}_{\lambda}^{\mathrm{mot}}\to\hat{\Pi}_{\lambda}\times_{\hat{\mathbb{Z}}}\mathrm{Hom}(\mathcal{W}_{p},\mathbb{G}_{m})$
are both isomorphisms.
Remark 2.3.
Ultimately these results (and [8, Proposition 3.5.2] rely on class field theory and its bootstrapping to the semisimplicity of geometric monodromy.
  1. (1)
    Concretely, for $\lambda\nmid p$ the proof that a homomorphism $\rho_{\lambda}\colon\hat{\Pi}_{\lambda}\times_{\hat{\mathbb{Z}}}\mathrm{Hom}(\mathcal{P}_{p},\mathbb{G}_{m})\twoheadrightarrow H$ classifies a $\lambda$-adic representation that is plain of characteristic $p$ (and similarly for the “weakly motivic” $\mathcal{W}_{p}$ case) can be interpreted as follows: the plain (or weakly motivic) property can be checked after restriction to a finite-index subgroup of $\Pi$, and from the semisimplicity of geometric monodromy one can check that after replacing $X$ by a suitable finite étale connected cover, $\rho_{\lambda}$ can be written as a product of homomorphisms $\hat{\Pi}_{\lambda}\to H^{\mathrm{geo},0}$ and $\mathrm{Hom}(\mathcal{P}_{p},\mathbb{G}_{m})\to Z_{H^{0}}^{0}$, with $H^{\mathrm{geo}}$ denoting the geometric monodromy group. The former is by Lafforgue’s theorem pure weight zero in any finite-dimensional representation, and the latter corresponds (in a choice of basis of $Z_{H^{0}}^{0}\cong\mathbb{G}_{m}^{r}$) to an $r$-tuple of elements of $\mathcal{P}_{p}$ that make the plainness conclusion evident. The converse (that plain representations of $\Pi$ correspond to homomorphisms out of $\hat{\Pi}_{\lambda}^{\mathrm{red}}$ that factor through $\hat{\Pi}_{\lambda}\times_{\hat{\mathbb{Z}}}\mathrm{Hom}(\mathcal{P}_{p},\mathbb{G}_{m})$) can be argued similarly.
  2. (2)
    To show a homomorphism $\Pi\to H(\overline{\mathbb{Q}}_{\lambda})$ is plain (respectively, weakly motivic) of characteristic $p$, it suffices to know $\rho_{\lambda}(\mathrm{Frob}_{x})$ has plain (or #$\#\kappa(x)$-Weil numbers, of possibly different weights) in a faithful representation for a single closed point $x$: compare [5, Proposition 4.3]. This can be seen from the sketch in item (1), or in terms of our proof by considering the restriction of $\rho_{\lambda}$ to the subgroup $1\times\mathrm{Hom}(\mathcal{P}_{p}/\mu_{\infty},\mathbb{G}_{m})\xrightarrow{\rho_{\lambda}}Z_{H^{0}}^{0}$: up to roots of unity, this homomorphism is determined by what it does on a single Frobenius element.
  3. (3)
    Since $\mathcal{P}_{p}/\mu_{\infty}$ is a $\mathbb{Q}$-vector space, the “plain” case of the Lemma can be deduced directly from the criterion of [8, Lemma 3.5.1], as Drinfeld does for $\hat{\Pi}_{\lambda}^{\mathrm{red}}$. Note that $\mathcal{W}_{p}/\mu_{\infty}$ is not a $\mathbb{Q}$-vector space, so in this case arguing as in Lemma 2.2 seems to be necessary.
Drinfeld’s main theorem ([8, Theorem 1.4.1], and see [10, Theorem 3.8] for the details when $\lambda|p$) gives for all $\lambda$ and $\lambda^{\prime}$ a unique isomorphism $\hat{\Pi}_{(\lambda)}\simeq\hat{\Pi}_{(\lambda^{\prime})}$ in $\operatorname{Pro-ss}(\overline{\mathbb{Q}})$ (the category of pro-semisimple groups, with homomorphisms taken up to conjugation by the identity component) preserving the canonical diagrams of Equation (1). He thus obtains an object $\hat{\Pi}$ of $\operatorname{Pro-ss}(\overline{\mathbb{Q}})$ that is unique up to a unique isomorphism.
Corollary 2.4.
For all $\lambda$ and $\lambda^{\prime}$ there is a unique isomorphism $\hat{\Pi}_{(\lambda)}^{\mathrm{plain}}\xrightarrow{\sim}\hat{\Pi}_{(\lambda^{\prime})}^{\mathrm{plain}}$ sending the canonical diagram
$\Pi_{\mathrm{Frob}}\to[\hat{\Pi}_{(\lambda)}^{\mathrm{plain}}{\,\!\sslash\!\,}\hat{\Pi}_{(\lambda)}^{\mathrm{plain},0}](\overline{\mathbb{Q}})\to\Pi$ (2)
to the corresponding diagram with $\lambda^{\prime}$ in place of $\lambda$. Similarly, there is a unique such isomorphism $\hat{\Pi}_{(\lambda)}^{\mathrm{mot}}\xrightarrow{\sim}\hat{\Pi}_{(\lambda^{\prime})}^{\mathrm{mot}}$.
Proof.
Under the isomorphism of Lemma 2.2, the isomorphism $\hat{\Pi}^{\mathrm{plain}}_{(\lambda)}\xrightarrow{\sim}\hat{\Pi}^{\mathrm{plain}}_{(\lambda^{\prime})}$ is induced by the unique isomorphism $\hat{\Pi}_{(\lambda)}\xrightarrow{\sim}\hat{\Pi}_{(\lambda^{\prime})}$ preserving Equation (1) and the identity on $\mathrm{Hom}(\mathcal{P}_{p},\mathbb{G}_{m})$. For the uniqueness, it suffices (by the Čebotarev density theorem and the uniqueness in [8, Theorem 1.4.1]) to check that an (on the nose) automorphism $\alpha$ of $\hat{\Pi}^{\mathrm{plain}}_{\lambda}$ that is trivial on $[\hat{\Pi}_{\lambda}^{\mathrm{plain}}{\,\!\sslash\!\,}\hat{\Pi}_{\lambda}^{\mathrm{plain},0}]$ and induces the identity on the pro-semisimple quotient $\hat{\Pi}_{\lambda}$ is the identity. Since $\alpha$ induces the identity on $\hat{\Pi}_{\lambda}$, it has the form $\alpha=\mathrm{id}\cdot\chi$ for some homomorphism $\chi\colon\hat{\Pi}_{\lambda}^{\mathrm{plain}}\to Z^{\mathrm{plain},0}\cong\mathrm{Hom}(\mathcal{P}_{p}/\mu_{\infty},\mathbb{G}_{m})$. Since
$[\hat{\Pi}_{\lambda}^{\mathrm{plain}}{\,\!\sslash\!\,}\hat{\Pi}_{\lambda}^{\mathrm{plain},0}]\cong[\hat{\Pi}_{\lambda}{\,\!\sslash\!\,}\hat{\Pi}_{\lambda}^{0}]\times_{\hat{\mathbb{Z}}}\mathrm{Hom}(\mathcal{P}_{p},\mathbb{G}_{m}),$
and $Z^{\mathrm{plain},0}$ injects into the right hand side (as $\ker(\mathrm{Hom}(\mathcal{P}_{p},\mathbb{G}_{m})\to\hat{\mathbb{Z}})$), we must have $\chi=1$, as needed. ∎
As in [8, §6], we can then define groups over $\overline{\mathbb{Q}}$, $\hat{\Pi}^{\mathrm{plain}}=\hat{\Pi}\times_{\hat{\mathbb{Z}}}\mathrm{Hom}(\mathcal{P}_{p},\mathbb{G}_{m})$ and $\hat{\Pi}^{\mathrm{mot}}=\hat{\Pi}\times_{\hat{\mathbb{Z}}}\mathrm{Hom}(\mathcal{W}_{p},\mathbb{G}_{m})$: these are objects of $\operatorname{Pro-red}(\overline{\mathbb{Q}})$ defined up to unique isomorphism, and for all $\lambda$, $\hat{\Pi}^{\mathrm{plain}}\times_{\overline{\mathbb{Q}}}\overline{\mathbb{Q}}_{\lambda}$ is isomorphic, uniquely up to conjugation by $\hat{\Pi}_{\lambda}^{\mathrm{plain},0}$, to $\hat{\Pi}_{\lambda}^{\mathrm{plain}}$ (and likewise for $\hat{\Pi}^{\mathrm{mot}}$).
We will apply the above conclusions in the following simplified form:
Corollary 2.5.
Let $\rho_{\lambda}\colon\hat{\Pi}_{\lambda}^{\mathrm{plain}}\to H_{\overline{\mathbb{Q}}_{\lambda}}$ be a surjection onto a reductive (not necessarily connected) group $H/\overline{\mathbb{Q}}$. Then $\rho_{\lambda}$ has a strong $\lambda^{\prime}$-companion $\rho_{\lambda\leadsto\lambda^{\prime}}$, also plain, for any finite place $\lambda^{\prime}$ of $\overline{\mathbb{Q}}$. In particular:
  • $H$ is the image of the homomorphism $\hat{\Pi}^{\mathrm{plain}}_{\lambda^{\prime}}\to H_{\overline{\mathbb{Q}}_{\lambda^{\prime}}}$ classifying $\rho_{\lambda\leadsto\lambda^{\prime}}$
  • For any closed point $x\in X$, the associated conjugacy class $\rho_{\lambda}(\mathrm{Frob}_{x})\in[H{\,\!\sslash\!\,}H](\overline{\mathbb{Q}}_{\lambda})$ is $\overline{\mathbb{Q}}$-rational and equal to the associated conjugacy class $\rho_{\lambda\leadsto\lambda^{\prime}}(\mathrm{Frob}_{x})$.
Here we write $\rho_{\lambda}(\mathrm{Frob}_{x})$ for the conjugacy class of the $\kappa(x)$-linearized Frobenius associated to the closed point, whether we are in the isocrystal or the lisse sheaf situation.
Proof.
@@
This is immediate from Corollary 2.4: descending $\rho_{\lambda}$ to $\overline{\mathbb{Q}}$, we get $\hat{\Pi}^{\mathrm{plain}}\twoheadrightarrow H$ (unique up to $H^{0}$-conjugation), which in turn for any $\lambda^{\prime}$ extends to the homomorphism $\hat{\Pi}_{\lambda^{\prime}}^{\mathrm{plain}}\twoheadrightarrow H_{\overline{\mathbb{Q}}_{\lambda^{\prime}}}$ classifying the companion $\rho_{\lambda\leadsto\lambda^{\prime}}$. ∎
Remark 2.6.
Note we do not claim the conjugacy class $\rho_{\lambda}(\mathrm{Frob}_{x})$ is independent of $\lambda$ as an element of $[H{\,\!\sslash\!\,}H^{0}](\overline{\mathbb{Q}})$ (compare the assertion in Corollary 2.4: indeed, this does not even make sense when only the closed point $x\in X$ rather than its lift $\tilde{x}\in\widetilde{X}$ is specified. The coarser nature of Corollary 2.5 allows us to transfer its conclusions to $\pi_{1}(X)$ without any ambiguity resulting from the non-canonical isomorphism $\pi_{1}(X)\cong\Pi$.